OpenStudy (tumblewolf):

The length of a rectangle is 3 times longer than its width. If the area of the rectangle is 24 square inches, then what is the perimeter of the rectangle? I know that the area formula is a=lw and the parameter is p=2l+2w. I just can't figure out what I'm supposed to do from here

Directrix (directrix):

 4W*W = 24 W^2 = 6 W = √6 L = 4√6 ---------- P = 2L + 2W p+= 10√6 P = 24.49 approx; 24.5 to the nearest tenth Solution courtesy of @Alan3354 @Tumblewolf